Show that T(x1, x2, x3, x4) = 3x1 −7x2 + 5x4 is a linear transformation by finding the

matrix for the transformation. Then find a basis for the null space of the transforma￾tion.


1
Expert's answer
2022-02-21T16:05:25-0500

T(x1, x2, x3, x4) = 3x1 −7x2 + 5x4 can be expressed as

"T\begin{bmatrix}
 x1 \\
 x2\\
x3\\
x4
\end{bmatrix} = \begin{bmatrix}
 3x1 \\
-7 x2\\
5x4\\
\end{bmatrix}"

T: R4"\to" R3

So the domain of T is R3. To find the column of the standard matrix of transformation we find T(e1), T(e2), T(e3), and T(e4)

Using the rule for transformation

T(e1) = "\begin{bmatrix}
 1 \\
0\\
0\\
0
\end{bmatrix} = \begin{bmatrix}
 3(1) \\
-7(0)\\
5(0)\\

\end{bmatrix} =
\begin{bmatrix}
 3 \\
0\\
0\\

\end{bmatrix}"


T(e2) = "\begin{bmatrix}
 0 \\
1\\
0\\
0
\end{bmatrix} = \begin{bmatrix}
 3(0) \\
-7(1)\\
5(0)\\

\end{bmatrix} =
\begin{bmatrix}
 0 \\
-7\\
0\\

\end{bmatrix}"


T(e3) = "\begin{bmatrix}
 0 \\
0\\
1\\
0
\end{bmatrix} = \begin{bmatrix}
 3(0) \\
-7(0)\\
5(0)\\

\end{bmatrix} =
\begin{bmatrix}
 0 \\
0\\
0\\

\end{bmatrix}"


T(e4) = "\begin{bmatrix}
 0 \\
0\\
0\\
1
\end{bmatrix} = \begin{bmatrix}
 3(0) \\
-7(0)\\
5(1)\\

\end{bmatrix} =
\begin{bmatrix}
 0 \\
0\\
5\\

\end{bmatrix}"


Therefore the transformation matrix can be expressed as

"\begin{bmatrix}
 3 & 0&0&0 \\
0&-7&0&0\\
0&0&0&5\\

\end{bmatrix}"


Finding the Basis for the null space of the transformation

"\begin{bmatrix}
 3 & 0&0&0 \\
0&-7&0&0\\
0&0&0&5\\

\end{bmatrix}

\begin{bmatrix}
 x1\\
x2\\
x3\\
x4

\end{bmatrix} = \begin{bmatrix}
 0 \\
0\\
0\\

\end{bmatrix}"


Reduced Row Echelon Matrix for the augmented matrix


"\begin{bmatrix}

 3 & 0 & 0 & 0 &\bigm| & 0 \\

 0 & -7 & 0 & 0 &\bigm| & 0 \\

 0 & 0 & 0 & 0 &\bigm| & 0 \\ 

 0 & 0 & 0 & 5 &\bigm| & 0

\end{bmatrix}"


Multiply row 1 by 1/3:

"\begin{bmatrix}

 1 & 0 & 0 & 0 &\bigm| & 0 \\

 0 & -7 & 0 & 0 &\bigm| & 0 \\

 0 & 0 & 0 & 0 &\bigm| & 0 \\ 

 0 & 0 & 0 & 5 &\bigm| & 0

\end{bmatrix}"


Multiply row 2 by -1/7:

"\begin{bmatrix}

 1 & 0 & 0 & 0 &\bigm| & 0 \\

 0 & 1 & 0 & 0 &\bigm| & 0 \\

 0 & 0 & 0 & 0 &\bigm| & 0 \\ 

 0 & 0 & 0 & 5 &\bigm| & 0

\end{bmatrix}"


Multiply row 3 by 1/5:

"\begin{bmatrix}

 1 & 0 & 0 & 0 &\bigm| & 0 \\

 0 & 1 & 0 & 0 &\bigm| & 0 \\

 0 & 0 & 0 & 0 &\bigm| & 0 \\ 

 0 & 0 & 0 & 1 &\bigm| & 0

\end{bmatrix}"


Convert the matrix equation back to an equivalent system:

"\begin{array}{rcrcrcr}
 x_{1}&&&&&=&\phantom{-}0 \\
 &&x_{2}&&&=&\phantom{-}0 \\
 &&&&x_{4}&=&\phantom{-}0 \\
 \end{array}"


Add an equation for each free variable:

"\begin{array}{rcrcrcrcr}
 x_{1}&&&&&&&=&\phantom{-}0 \\
 &&x_{2}&&&&&=&\phantom{-}0 \\
 &&&&x_{3}&&&=&x_{3} \\
 &&&&&&x_{4}&=&\phantom{-}0 \\
 \end{array}"

Solve for each variable in terms of the free variables:


"\begin{array}{rcrcrcrcr}
 x_{1}&&&&&&&=&\phantom{-}0 \\
 &&x_{2}&&&&&=&\phantom{-}0 \\
 &&&&x_{3}&&&=&x_{3} \\
 &&&&&&x_{4}&=&\phantom{-}0 \\
 \end{array}"


Collect terms into vectors:

"\left[ \begin{array}{c}x_{1}\\x_{2}\\x_{3}\\x_{4}\\\end{array} \right]
 = 
\left[ \begin{array}{c} 0\\0\\x_{3}\\0\\\end{array} \right]"


Factor out variables on the right side:


"\left[ \begin{array}{c}x_{1}\\x_{2}\\x_{3}\\x_{4}\\\end{array} \right]
 = 
x_{3}\left[ \begin{array}{c} \phantom{-}0\\\phantom{-}0\\\phantom{-}1\\\phantom{-}0\\\end{array} \right]"


Thus, a basis for the null space is:

"\displaystyle \left\{\left[ \begin{array}{c}
\phantom{-}0\\
\phantom{-}0\\
\phantom{-}1\\
\phantom{-}0\\
\end{array} \right]\right\}" [Answer]